A Carnot engine, with its cold body at 17°C has 50% efficiency. If the temperature of its hot body is now increased by 145°C, the efficiency becomes A. 55% B. 60% C. 40% D. 45%

Grade:upto college level

1 Answers

PRAKASHAM SETTY MANASANI
5 years ago
1case:TE=290k ,n=50%=1/2,TE=?      n=1-T2/TE=
1/2=1-290/TE
So T1=580k
Case
Now T1'=T1+418=580+418=998k
Now n'=1-T2/T1'==1-290/998=708/998
n'=708/998×100=70% nearly